A factory’s worker productivity is normally distributed.
One worker produces an average of 75 units
per day with a standard deviation of 20. Another
worker produces at an average rate of 65 per day
with a standard deviation of 21. What is the probability
that in 1 week (5 working days), worker 1
will outproduce worker 2?
1
Expert's answer
2020-12-10T11:55:26-0500

We have that

"\mu_1=75, \ \sigma_1=20"

"\mu_2=65, \ \sigma_2=21"

"n = 5"

Let's consider that "X_1" and "X_2" represent the production by worker 1 and worker 2 respectively. We need to find the probability that in 5 working days worker 1 will outproduce worker 2.

That is "P(\bar X_1>\bar X_2)"

"\bar X_1-\bar X_2" follows normal distribution with the mean "\mu=\mu_1-\mu_2=75-65=10"

and the standard deviation "\sigma=\sqrt{\frac{\sigma_1^2}{n}+\frac{\sigma_2^2}{n}}=\sqrt{\frac{20^2}{5}+\frac{21^2}{5}}=12.97"

"P(\bar X_1>\bar X_2)=P(\bar X_1-\bar X_2>0)=P(Z>\frac{0-\mu}{\sigma})=P(Z>\frac{-10}{12.97})="

"=P(Z>-0.77)=1-P(Z<-0.77)=1-0.2206=0.7794"


Answer: the probability that in 5 working days worker1 will outproduce worker2 is 0.7794.